OpenStudy (hmarie29):

the length of a rectangle is 2 inches more than three times the width .if the perimeter is 28 inches , what is the length of the rectangle? let l= length and w = width

OpenStudy (ben00):

Try dividing?

OpenStudy (hmarie29):

its telling me to show it in standerd form

OpenStudy (hmarie29):

such as l = w + 2 2l + 2w = 28

OpenStudy (ben00):

21 + 2w = 28 looks easy.

OpenStudy (hmarie29):

its 2L not 21

OpenStudy (ben00):

oh

OpenStudy (hmarie29):

yeah